OpenStudy (anonymous):

METAL FOR HELP Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% Using the function A(t)=P(1+rn)nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down. Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal.

OpenStudy (anonymous):

i have the first parts now i need to do this: Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+rn)1c]cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+rn.

OpenStudy (anonymous):

1.\[A(t)=39,145\left( \frac{ 1+0.03 }{ 12 } \right)^{12t}\] 2.\[\frac{ \log_{\frac{1+0.03}{12}}\left[\frac{ 10,000 }{ 39,145 } \right] }{ 12 }=t\]

OpenStudy (anonymous):

thats what i have so far

OpenStudy (anonymous):

Look up a calculator to do the job. If you're lucky the calculator you find explains itself.

OpenStudy (anonymous):

what do you mean a calculator

OpenStudy (anonymous):

If you know the type of problem you are doing, then type in on google: [type of problem] calculator, you'll probably get one that can answer your question.

OpenStudy (anonymous):

They help when you have a tough problem.

OpenStudy (anonymous):

okay i will try

OpenStudy (anonymous):

@jdoe0001 will u help me

undeadknight26 (undeadknight26):

What do you need now o.o Have you not already solved it?

OpenStudy (anonymous):

no Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+rn)1c]cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+rn.

OpenStudy (anonymous):

im really bad at word problems and logs and math

undeadknight26 (undeadknight26):

Im only in 9th and im in summer party mode...so taking me a lil while to get back into school mode XD

OpenStudy (anonymous):

crap bro im in 11th

undeadknight26 (undeadknight26):

No problem ive helped college students before...

undeadknight26 (undeadknight26):

I could solve this or i could simply google it and confirm it. This is your answer i believe: https://answers.yahoo.com/question/index?qid=20140727145333AAiNddT

undeadknight26 (undeadknight26):

btw its medal =.= Not metal :p

OpenStudy (anonymous):

this only shows one and two

OpenStudy (anonymous):

i need 3